# Optimizing Breast Cancer Detection in Mammograms: A Comprehensive Study of Transfer Learning, Resolution Reduction, and Multi-View Classification

### v2

@@ 3.6.1 Patch Classifier and Base-Model
Table 13: Classification accuracies of the patch classifiers evaluated on the VinDr-Mammo dataset.
@@
3.6.2 Patch Based Classifier (PBC)
Using the weights obtained from the patch classifier, we constructed single-view patch-based classification (PBC) models and achieved the results presented in Table 14. Training the ConvNeXt-Base (22k) model required approximately 24 hours for the three rounds on an NVIDIA A100 GPU with 40 GB of memory, making it the most time-intensive single-view model in this study.
Table 14: AUCs of the single-view classifier based on patches (PBC) for VinDr-Mammo.
@@ 3.6.2 Patch Based Classifier (PBC)
3.6.3 Direct Classifier (DC)
The results of the Direct Classifier (DC) experiments are summarized in Table 15. DenseNet169 emerged as the top-performing base model. Notably, the best-performing Patch-Based Classifier (PBC), ConvNeXt_Base_22k, showed a significant drop in performance when applied to the DC approach. Conversely, DenseNet169, which excelled in the DC task, performed poorly in the PBC setting.
Table 15: AUCs of the direct single-view classifier (DC) for VinDr-Mammo.
@@ 3.6.3 Direct Classifier (DC)
3.6.4 Conclusions about PBC
We compared the best PBC result (ConvNeXt_Base_22k, 0.8510±0.0163, Table 14) with the best DC result (DenseNet169, 0.8134±0.0136, Table 15) by running the DeLong test and obtained the value $p$ = 0.0013. This indicates with statistical significance that the best PBC model is superior to the best DC model for classifying high-quality mammograms. Due to this result, we recommend the PBC approach for the VinDr-Mammo set with 100% digital mammograms, despite the fact that it requires more training time.
On the other hand, four base models performed better in the Patch-Based Classifier (PBC) approach, while three models showed superior performance in the Direct Classifier (DC) approach (see Fig. 4 and Tables 14 and 15). Based on these results, it is not possible to conclude that the PBC approach is universally superior to the DC approach for classifying high-quality mammograms, as performance varies depending on the base model.
3.6.5 Conclusions on Base Models
3.7 Classifiers with Resized Images in VinDr-Mammo
To answer the question (3) for VinDr-Mammo, we downsample mammograms to half their size before classifying them.
3.7.1 Fixed Resizing Classifier (FRC)
After resizing the images by half using OpenCV’s INTER_AREA interpolation, we obtained the classification results presented in Table 16.
Table 16: AUCs of Fixed Resizing Classifiers (FRCs) on the VinDr-Mammo dataset.
@@ 3.7.1 Fixed Resizing Classifier (FRC)
3.7.2 Learn-to-Resize Classifier (LRC)
After reducing the resolution of the mammograms by half using the “learn to resize” technique, we obtained the classification results presented in Table 17.
Table 17: AUCs of Learn-to-Resize Classifiers (LRCs) on the VinDr-Mammo dataset.
@@ 3.7.2 Learn-to-Resize Classifier (LRC)
3.7.3 Conclusions on Resizing
As in Section 3.3.3, the top-performing Learn-to-Resize Classifier (LRC) demonstrated inferior performance compared to the best Fixed Resizing Classifier (FRC). Since LRC is both computationally more complex and resource-intensive than FRC, the Learn-to-Resize approach offers no practical benefit.
3.8 Two-View Classifiers in VinDr-Mammo
To address question (4), we followed the same network creation and training process outlined in previous sectionshae. We constructed a two-view classifier using the ConvNeXt_Base_22k network, initializing it with weights from the single-view PBC approach. Additionally, we processed each view (CC and MLO) individually using the best-performing classifier from the previous section (ConvNeXt_Base_22k) and combined the results using both average and maximum operations. The results are shown in Table 18.
Model Two-View Average Maximum
[…]
4 Conclusions
Reducing mammogram resolution consistently diminished model performance, especially for high-quality digital images. Furthermore, employing a two-view classification approach substantially outperformed single-view classifiers across both datasets (CBIS-DDSM: p-values under 0.0286; VinDr-Mammo: p-values under 0.0030). This strongly supports the adoption of multi-view strategies in future mammogram analysis frameworks.

### v3

@@ 4 Conclusions
Table S6: Classification accuracies of the patch classifiers evaluated on the VinDr-Mammo dataset.
@@
Table S7. Patch Based Classifier (PBC) (VinDr-Mammo) results
Table S7: AUCs of the single-view classifier based on patches (PBC) for VinDr-Mammo.
@@
Table S8. Direct Classifier (DC) (VinDr-Mammo) results
Table S8: AUCs of the direct single-view classifier (DC) for VinDr-Mammo.
@@
Table S9. Fixed Resizing Classifier (FRC) (VinDr-Mammo) results
Table S9: AUCs of Fixed Resizing Classifiers (FRCs) on the VinDr-Mammo dataset.
@@
Table S10. Learn-to-Resize Classifier (LRC) (FRC) (VinDr-Mammo) results
Table S10: AUCs of Learn-to-Resize Classifiers (LRCs) on the VinDr-Mammo dataset.